OpenStudy (anonymous):

Find the distance between the points (-1, 3) and (5, 0)

OpenStudy (anonymous):

(5+1)square+3 square =45

OpenStudy (anonymous):

(6,3)

OpenStudy (anonymous):

@ccalle ? are you there?

OpenStudy (texaschic101):

d = sq rt of (x2 - x1)^2 + (y2 - y1)^2 (-1,3) x1 = -1, y1 = 3 (5,0) x2 = 5, y2 = 0 lets sub now... d = sq rt (5 - (-1))^2 + (0 - 3)^2 d = sq rt (5 + 1)^2 + (-3)^2 d = sq rt 6^2 - 3^2 d = sq rt 36 - 9 d = sq rt 27 d = 5.196